(5%) Problem 6: The lead in pencils is a graphite composition with a Youngs modulus of about 1.10 N/m. AL Calculate how mu
Answer #1

Area is given by

A = pi*d2/4 = pi*(0.51*10-3)2/4

A=2.043*10-7 m2

Young's Modulus is given by

FL Y =AAL

FL 3.2 x 0.063 AL= AY - (2.043 x 10-7)(1 x 109) 3.

AL = 0.987mm